RC Meeting, Saturday, Dec 11, 2021
RC Meeting Information
Have RC meeting today
RC Daily Meetings are at 10 AM, remote
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-18.104.22.168 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
- Fall 2021 Shift Schedule
- PionLT Physics Status
- PionLT Runplans
- Hall-C Controls
- Hall-C Logbook
- MCC White Board
- OPS 7:45 RC meeting bluejeans link, weekends usually 9am.
- OPS daily 8:00 meeting and Wed 1:30 bluejeans link
Current RC: Salina Ali
Day shift: Adjusted spectrometer angles to prepare for HEEP singles
- Not taking data currently, because of magnet issue. Beam was off earlier but was ready to send to hall at approx. 9:00 am.
- Q3 on SHMS is down, Amy Comer is magnet on-call. She has rebooted it (9:30 AM).
- Power glitches throughout the accelerator seem to be the cause of the magnet issue
- Q3 is currently ramping up. Will start HEEP singles plan after Q3 is stable.
- Beam is ready to go, ready for HEEP singles
- Possible target issue: low pressure ~18-16 psi for LH2 target. Usually the acceptable pressure is 22 psi. Would become a major issue if it gets below atmospheric pressure (14.7 psi). Dave Meekins might come in later to find a way to increase pressure.
- The pressure was around the same for LH2 when running on LD2 in previous run plan
Update on heep singles run plan (8 GeV) day shift
- 11:04 AM: First runs with LH2 (10 mins) and Al dummy (6 min) at 30 uA was successful. HMS at 19.88 deg and SHMS at 8.26 deg
- 11:21 AM: When moving to next setting (HMS from 19.88 deg to 23.52 deg) and HMS got stuck at 20.7 degrees. There is a drive controller error [See log entry here]. Amy Comer is coming in to reset the circuit breaker and do a visual inspection of the tracks. She suspects the reset might not solve the issue, so we are skipping to the next setting, HMS at 25.02 deg, SHMS at 10.26 deg, taking Al data at 70 uA for 6 min, LH2 at 70 uA for 10 min.
- 12:28 PM: Amy has arrived, she will be going in the hall to reset the circuit breaker and do a visual inspection.
- 12:47 PM: HMS rotation was successful after circuit breaker reset and visual inspection. HMS is at 25.02 deg, and SHMS was moved to 10.26 deg for next setting.
- 1:00 PM: Greg Smith called me and told me it is a very serious issue. Dave Meekins was informed and plans to go in the hall to investigate. RadCon was called to make arrangements for the controlled access. We do not know how long this will take.
- 1:30 pm: Dave Meekins noticed one of the main supply valves (connected to ballast tanks) going to the target chamber for LH2 was off, and he had to manually turn it on. It is not known what or who turned this valve off for LH2. The pressure is now stable for LH2, and we are ready to take data for LH2. The pressure limits for the alarm handler have also been adjusted to 20-22 psi (previously they were 14-17 psi..).
- 2:00 pm: Beam is down for 30 more minutes, according to MCC. We will start data taking for LH2 after we have beam permit.
- 3:00 pm-swing shift 5:30 PM: Beam is running, continuing with heep singles program. Running at 40 uA, cannot get to 70 uA due to limitations from MCC.
- 5:45 PM, no beam. Hall C is in controlled access. Raster expert is coming in to solve issue with raster pattern. On oscilloscope raster has a square rotating shape.
- 6:30 PM: William has fixed the issue with the raster pattern. Now we can require 45 uA per raster coil to adhere to 70 uA requirement for heep singles.
- After HEEP singles, do luminosity scans, and then EDTM run
- Photo cathode forecast
- Steve Lassiter confirmed that moving to larger angles on Saturday does NOT require remote oversight from Amy or him.
- Check with Steve about moving to 6.16 degrees on Dec 20th. This can be done with remote oversight.
- Need to be sure that the Gui angle is modified to agree with the true angle when at 8.16 deg.
- Tues Dec 14th, there is a possible OPS test plan for 9-1 and Hall access.
Current Run Plan
We are currently on 8.0 GeV Part 1 [See run plan]
- Day shift:
- Running heep singles. [See run plan]
- 30 uA: Setting #1 HMS at 19.88 deg, PHMS=-5.2 GeV/c; SHMS at 8.26 deg, PSHMS=7.337 GeV on LH2 10 min, and Al dummy at 6 min. DONE
- Running heep singles. [See run plan]
Preparing for next setting, most likely (due to spectrometer rotation issues) we will be skipping the two settings for HMS at 23.53 deg and moving to HMS at 25.02 deg, SHMS at 10.26 deg
- 12:50 PM - Setting #2 Running at 40 uA: HMS at 25.01 deg, PHMS=-4.395 GeV/c; SHMS at 10.26 deg, PSHMS=-6.538 GeV on Al dummy for 6 min DONE
- 3:00 PM- Setting #2 Running 40 uA: HMS at 25.01 deg, PHMS=-4.395 GeV; SHMS at 10.26 deg, PSHMS=-6.538 GeV on LH2 for 10 min. DONE
- 4:45 PM: MCC can't give us more than 40 uA. Possible raster issue, but may not be an issue according to program deputy. Asking for 75 uA, but need to get approval from engineering to increase enabling current (Fig. 6) from 50 uA to 75 uA. Hopefully this will allow us to run at 70 uA. see log entry here
- 3:50 PM: Setting #3: Running 45 uA: HMS at 25.01 deg, PHMS=-4.395 GeV; SHMS at 17.79 deg, PSHMS=-5.512 GeV on LH2 for 10 min. SHMS Momentum changed right before swing shift. DONE
- 6:50: Setting #3: Waiting for beam. Running 40 uA: HMS at 25.01 deg, PHMS=-4.395 GeV; SHMS at 17.79 deg, PSHMS=-5.512 GeV on Al for 6 min.
- Raster issue is fixed. Ax and Ay coils are on, normal raster pattern
Notices in Effect for Shift Crew
- The Raster Coil pair Ax and Ay are off. Raster needs to be at 4x4.
- With LHe alarms it is good to check the cameras for plumes. If no plume and it is below the LO alarm and above LOLO then monitor with a plot and halog the problem.
- With LNitrogen alarms are not that critical. Just need to start monitoring the problem with plot and Halog.
- William Gunning came in around 6:30 PM and fixed the Raster issue. We had to plan an escorted access for this.
- There is a rotation issue that has come up if moving HMS between 19 to 25 degrees. It usually gets stuck around this range, please be wary of this issue. We will likely skip settings in this range if this drive controller error cannot be solved. For singles it will not be a big deal to skip settings for HMS at those angles.
- UPDATE 2:00 PM: Target pressure issue has been fixed
- Old: target pressure for loop 2, LH2, is lower than usual - around 17-18 psi. The alarm went off early in the day shift which brought it to the target operator's attention (Bill Henry). Usually acceptable pressure is around 22 psi. If this drops to atmospheric pressure, 14.7 psi, then we have a problem. The pressure for LH2 was around this range 17-18 psi during LD2 runs as well. We did contact target expert JP to resolve issue but it is currently ongoing and UNRESOLVED. We assume we can continue to take data. Most likely this would involve an access to the hall. See log entry from Greg Smith for an update on this issue (12:15 PM 12/11/21)
- The Raster Coil pair Ax and Ay are off
- Diffuser limit at 46uA
Crew Chiefs, We have permission to raise the Hall C diffuser raster limit to allow running >46uA to Hall C. This limit is normally imposed when one of the two rasters is not operative. Since the hall was able to double the running raster current to get the same size raster on the target as having both rasters, Keith is allowing us to raise the limit. To get to the setting, open the Hall C diffuser interlocks screen. Go to the bottom right corner and open the Expert screen. Just below and to the right of the FSD bits, raise the "Raster 1 OK Limit" to slightly above the highest current the hall wants to run at. Do not change the Raster 2 limit. Make a log entry with screen shots. Make an ABIL entry, System MPS and an expected removal date of 3 Jan. 2022. Paul
Planned Accesses to the Hall
- Maintenance to pump down HMS transfer line.
- Need to clean the SHMS encoder wheel
Planned Accelerator Configuration Changes